7+[(2+4)x2]-8+9x4=
please help me, i know rule is parenthesis first but then i'm confused because at end of problem there's multiplication and i thought would get done first

Answer by richwmiller(17219)

PEMDAS
parentheses exponents multiplication division addition subtraction
Start as far in as you can.
There are parentheses inside of brackets
7+[(2+4)*2]-8+9*4=
7+[6*2]-8+9*4=
7+12-8+9*4=
7+12-8+36=
55-8=
47
Try to avoid using x for multiplication.
When you get into algebra x is often used for variables.